Solve the following lengendre differential equation.

(1-x2)y"-2xy'+l(l+1)y=0


1
Expert's answer
2021-07-12T16:03:58-0400

Solution

(1-x2)y"-2xy'+l(l+1)y=0

Can be rewritten as;

y"-x2y"-2xy'+l(l+1)y=0

Solve using power series;

By definition;

y="\displaystyle\sum_{n=0}^
∞" anxn

And;

y'="\displaystyle\sum_{n=1}^
∞" nanxn-1

y"="\displaystyle\sum_{n=2}^
∞" n(n-1)anxn-2

Replace the values in the equation to obtain;

"\displaystyle\sum_{n=2}^∞" n(n-1)anxn-2-x2"\displaystyle\sum_{n=2}^∞" n(n-1)anxn-2-2x"\displaystyle\sum_{n=1}^∞" nanxn-1+l(l+1)"\displaystyle\sum_{n=0}^∞" anxn

Make all powers of X equal

"\displaystyle\sum_{n=0}^∞" (n+2)(n+1)an+2xn+"\displaystyle\sum_{n=2}^∞" -n(n-1)anxn+"\displaystyle\sum_{n=1}^∞" -2nanxn+"\displaystyle\sum_{n=0}^∞" l(l+1)anxn

n=0,1,2,3...

When;

n=0

((2.1)a2+l(l+1))x0=0

n=1

3.2a3-2a1+l(l+1)x1=0

n"=\geq2"

There is a general trend of

((n+2)(n+1)an+2-n(n-1)an-2nan+l(l+1)an))xn=0

All the coefficients of x must equal to zero, therefore;

Using

n=0;a2="\frac{-l(l+1)a_0}{2.1}"

n=1;a3="\frac{-(l^2+l-2)a_1}{3.2}"

n"\geq" 2

an+2="\frac{(n+l+1)(n-l)a_n}{(n+2)(n+1)}"

This is the recurrence relation.

Using it ;

n=2,

a4="\frac{(l+3)(2-l)a_2}{4.3}" ="\frac{(l+3)(l+1)(l)(l-2)a_0}{4!}"

n=3

a5="\frac{(l+4)(3-l)a_3}{5.4}" ="\frac{(l+4)(l+2)(l-1)(l)(l-3)a_1}{5!}"

And so on

A pattern for even number is ;

"a_{2n}=\frac{(-1)^n(l+2n-1)(l+2n-3)...(l+1)l(l-2)...(l-(2n-2))}{(2n)!}"

A pattern for the odd numbers is;

"a_{2n+1}=\frac{(-1)^n(l+2n)(l+2n-2)...(l+2)(l-1)(l-3)...(l-(2n-1))a_1}{(2_{n+1})!}"

Since y="\displaystyle\sum_{n=0}^∞" anxn=a0"\displaystyle\sum_{n=0}^∞" "\frac{(-1)^n(l+2n-1)(l+2n-3)...(l+1)l(l-2)..(l-(2n-2))}{(2n)!}x^{2n}+a_1\displaystyle\sum_{n=0}^∞\frac{((-1)^n(l+2n)(l+2n-2)..(l+2)(l-1)(l-3)...(l-(2n-1))}{(2n+1)!}x^{2n+1}"